One number is five more than a second number. The product of the two numbers is 374. Use guess and check or other means to find the two numbers
Answer by scott8148(6628)

"guess and check"???? __ this is ALGEBRA.com

let x="2nd number", so x+5="1st number"

"The product of the two numbers is 374" __ x(x+5)=374 __ x^2+5x-374=0

factoring __ (x+22)(x-17)=0 __ so x=-22 and x=17

the two numbers are 22 and 17 __ they are either both positive or both negative